There are two facing pages, The product of the page number is 1190. Find the page numbers of the first page and second page?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16949)

If one page is x
the other will be (x+1)
x(x+1)=1190
x^2+x=1190
x^2+x-1190
x^2+35x-34x-1190
x(x+35)-34(x+35)=0
(x+35)(x-34)
x= 34
The pages are 34,35

Answer by stanbon(75887)

There are two facing pages, The product of the page number is 1190. Find the page numbers of the first page and second page?
----
Let the 1st page number be "x".
The next is "x+1":
Equation:
x(x+1) = 1190
x^2+x-1190 = 0
(x+35)(x-34) = 0
Positive solution:
x = 34
x+1 = 35
============
Cheers,
Stan H.
==============